haz:

Verify the identity. cot(x-pi/2)=-tan x

supie:

@darkknight

haz:

\[\cot(x-\frac{ \pi }{ 2 })=-\tan x\]

darkknight:

are we trying to verify if this is a trig identity or not?

haz:

yes

darkknight:

ok im assuming you are in at least precalc, so use this https://calcworkshop.com/pdf/trig-identities-cheatsheet.pdf to refer to all of the basic trig identities, we use these to see if your identity is correct

haz:

i have some of those but i just dont know where to start

darkknight:

im in class rn, can you wait like an hour or so?

haz:

sure, thank you

darkknight:

np

jhonyy9:

i think will help you use \[\cot x = \frac{ \cos x }{ \sin x } and \tan x = \frac{ \sin x }{ \cos x }\]

jhonyy9:

or using this formula will help you Cot(A-B) = (CotACotB+1)/(CotB-CotA)

jhonyy9:

remember that Cotθ = 1/tanθ

jhonyy9:

and you need to know again that cot(pi/2) = 0